If the legs in a right triangle are 1 and root 2 what is the hypotenuse?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

legs are 1 and sqrt2
hyp^2= 1^2 +sqrt2^2
hyp^2= 1 +2
hyp^2= 3
hyp = sqrt 3
Ananth

Answer by richwmiller(17219)

1^2+(sqrt(2))^2=c^2
1+2=c^2
3=c^2
c=sqrt(3)